Reports 1-1 of 1 Clear search Modify search
AdV-TCS (CO2 laser projector)
berni, bersanetti, boldrini, lumaca - 0:22 Thursday 16 March 2023 (59274) Print this report
DAS power differential tuning

After the activities performed this morning on ALS (59266) and some last preliminary work for lock stability related to the glitches on PSTAB (59259), at around 19.00 UTC the activity for the DAS power differential tuning started.

The ITF has been locked at CARM NULL 1F since 17.20 UTC, at its nominal working point (59271).

The proposed actions on DAS actuators power were to:

  • reduce the WI DAS by -10%

  • increase the NI DAS by +10%

 

The starting point was:

0

 

CH [W]

INNER DAS [W]

OUTER DAS [W]

W

on the ITF

0.065

0.068

1.41

on the pickoff

0.399

0.011

0.230

N

on the ITF

0.079

0.215 + 0.08 (cold lens) = 0.295

1.35 + 1.8 (cold lens) = 3.15

on the pickoff

0.487

0.048 (0.035+0.013)

0.513 (0.22+0.293)

 

First step, performed between 19.00-19.10 UTC:

1

 

CH [W]

INNER DAS [W]

OUTER DAS [W]

W

on the ITF

0.065

0.061

1.27

on the pickoff

0.399

0.010

0.207

N

on the ITF

0.079

0.237 + 0.08 (cold lens) = 0.317

1.49 + 1.8 (cold lens) = 3.29

on the pickoff

0.487

0.052 (0.039+0.013)

0.536 (0.243+0.293)

At 20.55 UTC, (almost 2 hrs after the power change, thermal transient mostly completed), the ITF was unlocked on purpose, to repeat the locking acquisition and realign the optics: at 21.22 UTC CARM NULL 1F was reached again, without troubles.

In Fig. 1 the behavior of the main signals during the shift are shown. Not big effect was observed, but the 12 MHz seemed to raise up and the dark fringe stopped to brightened starting from the DAS change moment.

In Fig. 2, 3, 4 the PC images immediately before the DAS change, after the thermal transient before unlock and around 1 hour after the second lock are shown, respectively.

 

At the end of the shift, the automatic measurement of absorption has been started on NI TM. The automatic measurement started at 22.09 UTC. The HWS DET python acquisition failed to start (see Diego’s entry for details). Diego managed to start the acquisition manually at 23.08 UTC, then we left the automation to continue the measurement: it should end by itself around 05.00 LT in the morning.

Images attached to this report
Search Help
×

Warning

Error

The present report has been modified outside this window. Please check for its integrity in the main page.

Refreshing this page will move this report into drafts.

×